OpenStudy (anonymous):

−3t/1−4/p=6/s solve for "p" ....please explain how to do it....

OpenStudy (anonymous):

are you sure that is typed correctly?

OpenStudy (anonymous):

probably w/o the number 1

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

are you there?

jhonyy9 (jhonyy9):

-3t -4/p =6/s -3tsp -4s =6p -4s=6p+3tsp -4s=p(6+3ts) p= -4s/(6+3ts)